OpenStudy (anonymous):

if a coin is flipped 10 times what is the probability of at most 5 heads given atleast 1 head?

OpenStudy (cwtan):

may i use probability distribution to solve it?

OpenStudy (anonymous):

no solve it by conditional probability

Parth (parthkohli):

Give at least one head = Doesn't give all tails Give at least one head = 1 - gives all tails Give at least one head = 1 - 1/1024 = 1023/1024

OpenStudy (cwtan):

u havent solve for most 5 head

OpenStudy (anonymous):

oh solve it for this too plz

OpenStudy (cwtan):

wa it seem.......

OpenStudy (anonymous):

plz solve it

OpenStudy (cwtan):

my brain stucked.....take time wait a while

OpenStudy (anonymous):

ok

Parth (parthkohli):

You can do it yourself lol it's simple

OpenStudy (anonymous):

ok